An LRT car, 5 km distance from the Tayuman, takes 100 A over a 100 mm hard-drawn copper  trolley wire having a resistance of 0.270 ohms per km. The rail and ground return has a resistance  of 0.06 ohm per km. If the station voltage is 750 V, what is the voltage of the car? 




1
Expert's answer
2021-10-06T08:12:36-0400

The combined resistance of the wires is


"R=\rho_w d+\rho_gd."

The voltage drop across the wires:


"V_d=IR=Id(\rho_w+\rho_g)."

The voltage of the car:


"V=V_s-V_d=V_s-Id(\rho_w+\rho_g),\\
V=750-100·5(0.27+0.06)=585\text{ V}."